A tablet contains 238 mg of magnesium chloride (MgCl2 MW = 95.2 g/mol). How many milliequivalents of magnesium ions does the tablet contain? Assume complete disassociation. (Round your answer to the nearest whole number of milliequivalents. How is this worked out?

The answer is 5
Answer #1

ANSWER.

* Calculation of milliequivalence of magnesium ion in the tablet.

        Formula

                 mEq = mg * valence / molar mass.

     Here

           mg = 238 mg.

           Valence = 2 (Mg ion has a valence of 2 because it has a charge of +2).

           Molar mass = 95.2 g/mol.

   As per above formula

              mEq = 238 * 2 / 95.2

                       = 476 / 95.2

                       = 5 mEq.